Consider the following hypothesis test for the mean difference. (Give your answers correct to four decimal places.)
(a) Determine the p-value for Ho: μd = 0 and Ha: μd > 0, with n = 20 and t = 2.


(b) Determine the p-value for Ho: μd = 0 and Ha: μd ≠ 0, with n = 19 and t = -1.89.


(c) Determine the p-value for Ho: μd = 0 and Ha: μd < 0, with n = 25 and t = -2.6.


(d) Determine the p-value for Ho: μd = 0.75 and Ha: μd > 0.75, with n = 7 and t = 3.48.


Answer by stanbon(75887)

Consider the following hypothesis test for the mean difference. (Give your answers correct to four decimal places.)
(a) Determine the p-value for Ho: μd = 0 and Ha: μd > 0, with n = 20 and t = 2.
p-value = P(t > 2 when df = 19) = tcdf(2,100,19) = 0.03
----------------------------
(b) Determine the p-value for Ho: μd = 0 and Ha: μd ≠ 0, with n = 19 and t = -1.89.
p-value = 2*P(t < -1.89 when df=18) = 2*tcdf(-100,-1.89,19) = 0.0750
-----------------------
(c) Determine the p-value for Ho: μd = 0 and Ha: μd < 0, with n = 25 and t = -2.6.
p-value = p(t < -2.6 when df = 24) = tcdf(-100,-2.6,24) = 0.0079
-----------------------------
(d) Determine the p-value for Ho: μd = 0.75 and Ha: μd > 0.75, with n = 7 and t = 3.48.
p-value = P(t > 3.48 when df = 6) = tcdf(3.48,100,6) = 0.0066
-----------
Cheers,
Stan H.
------------